OpenStudy (anonymous):

Solve the formula for the variable y. z = 4x + 2y A. y = z/2 - 4x B. y = z – 2x C. y = z/2 -2x D. y=z/2 + 2x

OpenStudy (anonymous):

z = 4x + 2y 2y + 4x = z -4x -4x 2y = z - 4x divide both sides by 2 y = (z-4x)/2 or y = 0.5(z-4x)

OpenStudy (anonymous):

or y = (z/2) - 2x

OpenStudy (anonymous):

Ah okay thank you!